ananazzaa:

Larry and Peggy are making decisions about their bank accounts. Larry wants to deposit $350 as a principal amount, with an interest of 4% compounded quarterly. Peggy wants to deposit $350 as the principal amount, with an interest of 6% compounded monthly. Explain which method results in more money after 2 years. Show all work.

WickedHeart:

@ananazzaa What would you say the answer would be?

ananazzaa:

I dont understand this problem at all

WickedHeart:

@ananazzaa wrote:
I dont understand this problem at all
Okay so you will use the Formula \[A = 350(1 + 0.04/4)(4 * 2)\] \[A = 350(1 + 0.06/12)(4 * 2)? \]

WickedHeart:

Let's look at the first one. \[A=350 x (1+ (.04/4))^4\] Try entering it like this in the calculator

Sophieee:

Larry- principle = $350 interest = 4% time = 2 years Compound interest for quarterly: Rate = R/4 = 1% Time = 4n = 4x2 = 8 years

Sophieee:

Using formula

Sophieee:

Peggy: Principle = $350 Interest = 6% Time = 2 years

Sophieee:

Compound Interest for monthly Rate

WickedHeart:

Alright the girl got it, but so you know no direct answers Allowed @sophieee

Sophieee:

Kk

ananazzaa:

lollll

ananazzaa:

thankyou sm

ananazzaa:

my bad @wickedheart

WickedHeart:

@ananazzaa wrote:
my bad @wickedheart
Its cool just letting you know next time

Sophieee:

Yeah I didn't know either I'm sorry

WickedHeart:

@ananazzaa Closed when done please^^

WickedHeart:

@sophieee wrote:
Yeah I didn't know either I'm sorry
Its all good